OpenStudy (anonymous):

Find the slope and the y-intercept of the equation 4x + 5y = -9. MEDAL WILL BE GIVEN

OpenStudy (anonymous):

okay so how would you normally go about finding the slope of an equation?

OpenStudy (anonymous):

y=mx+b?

OpenStudy (anonymous):

Yes, that's right. The first step is to solve for y in terms of x. So 4x+5y=-9 5y=-4x-9 y=-4/5x-9/5

OpenStudy (anonymous):

now you have y=mx+b; y=-4/5x-9/5

OpenStudy (anonymous):

what's your m?

OpenStudy (anonymous):

-4/5?

OpenStudy (anonymous):

Yep you got it

OpenStudy (anonymous):

only thing left in the problem is to find the y intercept. Do you know how to do that?

OpenStudy (anonymous):

A line parallel to the A.Y int. Or B.X-Int has no defined slope, and a line parallel to the A.Y int. Or B.X-Int has a slope of zero? do you know that

OpenStudy (anonymous):

don't you still need to find the y intercept?

OpenStudy (anonymous):

it was -9/5

OpenStudy (anonymous):

oh good. okay you got it.

OpenStudy (anonymous):

well, when does a line have a slope of zero?

OpenStudy (anonymous):

undefined?

OpenStudy (anonymous):

let me rephrase the question. what is the slope of a horizontal line?

OpenStudy (anonymous):

0?

OpenStudy (anonymous):

you are right but why?

OpenStudy (anonymous):

I don't know parallel? No slope?

OpenStudy (anonymous):

a horizontal line has the same why everywhere along the line. the only thing that changes is the x.

OpenStudy (anonymous):

y*

OpenStudy (anonymous):

so by the slope formula: y2-y1/x2-x1=slope=0 because y2-y1=0

OpenStudy (anonymous):

y2-y1=0 because the y never changes on a horizontal line. so consider a vertical line where the y is always changing but the x is always the same

OpenStudy (anonymous):

by the slope formula: y2-y1/x2-x1=slope=undefined. this is because x2-x1=0 and therefore; y2-y1/0=undefined since you can never divide by a 0

OpenStudy (anonymous):

does that make sense

OpenStudy (anonymous):

not really , so does a line parallel to the y axis has no defined slope?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

and a line parallel to the x-axis has a slope of zero?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

If a line with a slope of -2 crosses the y-axis at (0, 3), what is the equation of the line? would it be y+2x=3?

OpenStudy (anonymous):

yes